tem 7 Part A A transverse wave on a string is given by the formula D(z, 0.15 sin(4.0z 43.0t). where D and z are in m and t is in s At t _ 7.0 x 10 2s. wnat is the transverse displacement of the point on the string where z-0.45m? Express your answer using two significant figures. D(0.45m, 7.0 × 10-2S)- Submit My Answers Give Up Part B At t-7.0 × 10 2s. what is the (transverse) velocity of the point on the string where Express your answer using two significant figures. 0.45m? lndicate he direction or velocity with the son of your answer (0.45m, 7.0× 10 29- or m/s Submit My Answers Give Up

 

A transverse wave on a string is given by the formula D(x,t) = 0.15 sin(4.0x - 43.0t). where D and z are in m and t is in s 

Part A

At t = 7.0 x 10 2s. what is the transverse displacement of the point on the string where z-0.45m?

 Express your answer using two significant figures. 

D(0.45m, 7.0 × 10-2S) = 


Part B 

At t = 7.0 × 10 2s. what is the (transverse) velocity of the point on the string where x = 0.45m? 
Answer #1

(A)

Displacememt on the string is

D(0.5m , 7*10-2 s) = 0.1*sin(3.6(0.5) - 41(7.4*10-2) )

= 0.1sin(-1.234)

= 0.1*-0.0215

= -0.022 m

(B)

the velocity of the string is ,

\deltaD (x,t) /\deltat = 0.10(-41)cos(3.6*x - 41*t)

\deltaD (0.45 , 7*10-2) /\deltat = -4.1 cos(3.6*0.45 - 4.1*7*10-2)

= -4.1cos(1.62 - 0.287)

= -4.09 m/s
Answer #2

43モ = -45 m Sın ( r. 8-3-o↓ D , IS = -- 、14 3 m d D- -6.45 S ( i-8-3.01) - 2 2 s mls